OpenStudy (bananas1234):

1. Multiply and simplify. a) (b + 2)(b + 2) b) (x + 3)(x – 3)

OpenStudy (lalipapa123):

(b+2)(b+2) =(b+2)^2 = [ (a+b)^2 = a^2 + 2ab + b^2 =here a=b and b= 2 substitute is the equation you will get it

OpenStudy (bananas1234):

Thank you, so for the b part how did you get 2? @lalipapa123

OpenStudy (anonymous):

thats just the formula to foil (a+b)(a+b) = a^2 + ... (b+ 2)(b+2)= b^2 + ...

OpenStudy (bananas1234):

so for the a question. (b + 2)(b + 2) = (b+2)^2 = a^2 + 2ab + b^2 what do i do next, if i did the above correct?

OpenStudy (lalipapa123):

the second one is like (a+b) (a-b)=(a^2-b^2)

OpenStudy (bananas1234):